A stone is thrown with a velocity of 8 ms^-1 into the air at an angle of 40° to the horizontal.
(Give your answer to 2 significant figures).
a) Calculate the vertical component of the velocity.
b) State the value of the vertical component of the velocity when the stone reaches its highest point. Ignore air resistance.
c) Use your answers to a and b to calculate the time the stone takes to reach its highest point.
d)Calculate the horizontal component of the velocity.
e)Use your answers to c and d to find the horizontal distance travelled by the stone as it climbs to its highest point.

Answer by Alan3354(69443)

A stone is thrown with a velocity of 8 ms^-1 into the air at an angle of 40° to the horizontal.
(Give your answer to 2 significant figures).
a) Calculate the vertical component of the velocity.
vert = 8*sin(40) m/sec
--------
b) State the value of the vertical component of the velocity when the stone reaches its highest point. Ignore air resistance.
Zero (0).
-----------
c) Use your answers to a and b to calculate the time the stone takes to reach its highest point.
Using 9.8 m/sec/sec for gravity:
v = at ---> t = v/a
t = 8*sin(40)/9.8 to apogee
---------
d)Calculate the horizontal component of the velocity.
horiz = 8*cos(40)
----------
e)Use your answers to c and d to find the horizontal distance travelled by the stone as it climbs to its highest point.
horiz = 8*cos(40)*2t meters